Dividing Partnership Net Loss

Leigh Meadows and Byron Leef formed a partnership in which the partnership agreement provided for salary allowances of $38,000 and $33,000, respectively.

Determine the division of a $19,000 net loss for the current year, assuming remaining income or losses are shared equally by the two partners. Use the minus sign to indicate any deductions or deficiencies.

Leigh Meadows Byron Leef Total
Salary Allowance $ $ $
Remainder $ $ $
Net Loss $ $ $
Answer #1
Leigh Meadows Byron Leef Total
Salary Allowance 38000 33000 71000
Remainder -45000 -45000 -90000
Net Loss -7000 -12000 -19000
Workings:
Remainder loss =71000+19000 = 90000
Loss division =90000*50%=45000